OpenStudy (anonymous):

a line passes through a point (-2,3).find an equation of this line if it is a horizontal

OpenStudy (anonymous):

horizontal line have the form y=b where b is a real number; if the line goes through the point (-2,3) then the line has equation y=3. every point on the horizontal line has y-coordinate 3, btw.

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

thanks dear

OpenStudy (anonymous):

wat's the answer when it has slope 3/2?

OpenStudy (anonymous):

y-3=(3/2)(x+2) y=(3/2)x+3+3 y=(3/2)x+6

OpenStudy (anonymous):

mandolino u didn't include the slope in ur calculation